OpenStudy (anonymous):

find f⋅g given that: f(x)=x^2-√x, g(x)=2x-√x

OpenStudy (johnweldon1993):

So we need to multiply this out... \[\large (x^2 - \sqrt{x}) \times (2x - \sqrt{x})\] Kind of like foil... multiply the first term in the first parenthesis by every term in the second parenthesis... Then do the same with the second term in the first parenthesis..what do you get?